OpenStudy (anonymous):

A certain species of tree grows an average of 0.5 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 800 centimeters tall

OpenStudy (anonymous):

shud be able to times that by the week.. o.5 times by the week number

OpenStudy (anonymous):

?

OpenStudy (anonymous):

its 52 weeks in a year. startin at 800 after 52 weeks how much have the tree grown

OpenStudy (anonymous):

goin by ur statement. if it grows 0.5cm a week for a year how much is that

OpenStudy (anonymous):

182? I really do not know that is why I asked for help.

OpenStudy (anonymous):

decimals we are dealin with right.. so its 52 weeks in a year, 0.5 x 52 equal 26.. but that's just for a year?? idk how long u have to go.. but if it grows for a year. startin at 800.. ?? answer will be??? say it lol

OpenStudy (anonymous):

you kno it .. I just don't wanna give u the answer.. cause this site is to help u learn it.. but I can give u lots of clues

OpenStudy (anonymous):

Laughing out loud... Hold on.

OpenStudy (anonymous):

It's difficult because I have to put it in this formula: tn=800+(n-1)0.5, then simplify it.

OpenStudy (anonymous):

wow u actually spelled it ;lol ok I see formula. does it tell u how long ?? like if it grew for 6months ex.

OpenStudy (anonymous):

I know, everyone questions me on that. Not really, that's the problem.

OpenStudy (anonymous):

By the way, the question just asks for an equation for the sequence that represents the weekly height of this tree in centimeters.

OpenStudy (anonymous):

yeah... smh well if it grow for a year its 26cm being 0.5 x 52 = 26 so lets try n=800+(n-1)0.5 n wud b 26 n=800+(26-1)0.5 (25)0.5 n=800+12.5 ???

OpenStudy (anonymous):

oh ok that right there shud work for you.. being that you complete it

OpenStudy (anonymous):

I don't know if I should question you, because it asked for a weekly growth, that part is confusing me.

OpenStudy (anonymous):

okay that's why I said 52 to weeks.. its on u to mark each one at 0.5 week 1 800.5cm week2 801cm etc.. u understand now?? I use multi. cuz it was faster

OpenStudy (anonymous):

week 52 826cm

OpenStudy (anonymous):

Okay, that makes more sense. Thank You.

OpenStudy (anonymous):

you prolly wont hav to go that high. cuz they will see where you going w it

OpenStudy (anonymous):

One more thing, when you were simplifying the equation, why did you put question marks. n=800+12.5 ???

OpenStudy (anonymous):

and your Welcome just tryna help ;)

OpenStudy (anonymous):

because I was solving the equation.. lol but u dnt need to solve it u just need the equation

OpenStudy (anonymous):

Okay. Once again Thank You :)